OpenStudy (anonymous):

Use the Pythagorean theorem to find the length of the hypotenuse with one side measuring 16 feet and the other measuring 7 feet. 21.34 feet 17.46 feet 14.39 feet 16.78 feet

OpenStudy (anonymous):

a² + b² = c² 16² + 7² = c² 256 + 49 = c² 305 = c² c = √305 = 17.46 feet